### v1

@@ Abstract
For decades, researchers have developed task-specific models to address scientific challenges across diverse disciplines. Recently, large language models (LLMs) have shown enormous capabilities in handling general tasks; however, these models encounter difficulties in addressing real-world scientific problems, particularly in domains involving large-scale numerical data analysis, such as experimental high energy physics. This limitation is primarily due to BPE tokenization’s inefficacy with numerical data. In this paper, we propose a task-agnostic architecture, BBT-Neutron, which employs a binary tokenization method to facilitate pretraining on a mixture of textual and large-scale numerical experimental data. The project code is available at https://github.com/supersymmetry-technologies/bbt-neutron. We demonstrate the application of BBT-Neutron to Jet Origin Identification (JoI), a critical categorization challenge in high-energy physics that distinguishes jets originating from various quarks or gluons. Our results indicate that BBT-Neutron achieves comparable performance to state-of-the-art task-specific JoI models. Furthermore, we examine the scaling behavior of BBT-Neutron’s performance with increasing data volume, suggesting the potential for BBT-Neutron to serve as a foundational model for particle physics data analysis, with possible extensions to a broad spectrum of scientific computing applications for Big Science experiments, industrial manufacturing and spacial computing.
@@ 1 Introduction
Scaling law plays a critical role in developing emergent abilities and in-context learning of large language models [1, 2, 3, 4]. Pretraining LLM on a sufficiently broad and heterogeneous corpus gives rise to in-context learning building upon the processes of absorbing skills within the repetitive sequences of subsets of data which usually come as natural language [5, 6]. The efficacy of in-context learning is directly proportional to the volume and variety of the data, as well as the computational power expended during the pretraining phase, adhering to a power-law distribution.[7]. The emergence of LLM, described by the ability to solve specific tasks jumping up abruptly when data and model size pass a certain threshold, is attributed to the scaling of pretraining.
@@
First, the mainstream number representation method, Byte Pair Encoding (BPE) tokenization, as adopted by GPT-3[5, 6], limits large language models’ (LLMs) performance in numerical processing and reasoning tasks. BPE tokenization—a simple data compression technique that iteratively replaces the most frequent byte pairs with a single, unused byte—can introduce ambiguity when tokenizing numbers, as it often fragments numerical values into arbitrary segments, thereby obscuring their original meaning. Additionally, BPE’s statistical approach may result in inconsistent tokenization of identical numbers across different contexts, further complicating downstream numerical tasks. Llama3 [17] attempts to address this by splitting numbers into indi-vidual tokens to improve consistency in number representation. However, even the most advanced number tokenization techniques cannot eliminate the fundamental issue: the quantitative relationships between individual numbers are distorted post-tokenization, impairing foundational models’ ability to perform basic comparisons and arithmetic operations [9, 18, 19]. This limitation has driven consensus on the necessity of advancing number tokenization methods to enable significant improvements in LLM applications for scientific tasks.
@@
Third, foundational models have been applied to scientific tasks in areas such as question answering[23, 24], quantum physics calculations [25], prime number distribution [26], partial differential equation (PDE) computation [27], and time series regression [28]. However, to date, LLMs developed for scientific applications are primarily trained on symbolic data, including text and scientific symbols such as DNA sequences and mathematical formulas. Numerical experimental data—particularly from large-scale scientific projects in fields like particle physics and astronomy—are typically excluded from these training corpora due to the absence of a general architecture capable of multimodal synthesis of text and experimental data. This limitation significantly hinders the capacity of general LLMs to address complex, real-world scientific problems.
@@
JoI is a newly developed idea for experimental particle physics that aims to distinguish jets originating from different quarks and gluons, which are the building blocks of our material world and are fundamental particles in the Standard Model of Particle Physics. Once quark or gluon is produced in high-energy collisions, such as the Large Hadron Collider (LHC) and future electron-positron Higgs factories, it would immediately give rise to a spray of particles—mostly hadrons—that travel in the same direction. This spray of particles is conventionally known as a jet, which is a critical object for physics measurements in collider experiments. Significant efforts have been devoted to identifying the origin of jets—specifically, determining which type of colored particle initiated the jet. Identification of jet origin is central to many physics analyses, particularly in studies of the Higgs, W, and Z bosons, as nearly 70% of these bosons decay into two jets directly. In addition, jets are fundamental to advancing our understanding of QCD. However, jets originating from different types of colored particles exhibit only marginal differences in their observables, making accurately identifying the origin of the jet extremely challenging.
@@
JoI was incarnated in the specific algorithms of jet flavor tagging, jet charge measurement, and strange, light, and gluon jet tagging. At the Large Electron-Positron Collider (LEP), the predecessor to the LHC, jet origin identification initially relied on cuts applied to various discriminating variables to distinguish between b-, c-, and light-jets [30, 31]. Over time, likelihood-based methods [32, 33] emerged, combining multiple observables into a single discriminant. With advancements in detector technology, the LHC provided improved reconstruction performance. Coupled with advanced algorithms such as Boosted Decision Trees (BDTs) [34, 35], followed by recurrent neural networks (RNNs) [36, 37], convolutional neural networks (CNNs) [38, 39], Graph Neural Network (GNN), and transformer, the origin of jet been identified with better performance. JoI was then developed using GNN-based ParticleNet and Transfomer-based Particle Transformer; it could simultaneously identify 11 jet species—five quarks, five anti-quarks, and gluons—at the proposed electron-positron Higgs factory. JoI offers an unprecedented distinguishing power of different jet species and is critical for almost all the physics measurements with the jet’s final state. Quantitative analysis shows that the anticipated accuracies of relevant Higgs measurements could be improved from two-folds up to two orders of magnitude using JoI, and thus significantly boost the discovery power of collider experiments.
@@
We observe that BBT-Neutron achieves a performance comparable to that of specialized models like ParticleNet and Particle Transformer. As the first task-agnostic general LLM architecture capable of pretraining on a multimodal mixture of textual and large-scale numerical datasets, BBT-Neutron demonstrates state-of-the-art performance in JoI, compared to the results of leading domain-specific models. This achievement marks a significant step toward developing foundational models for scientific research applications that can generalize across diverse data types and tasks.
@@ 2 Results
We conduct a comparative analysis on scaling behaviors of BBT-Neutron, ParticleNet, and Particle Transformer applied to JOI across training datasets ranging from 100 to 10 million events.The results for these models are shown in the bottom panels of Fig. 2. $M_{11}$, which classifies each jet according to the category with the highest predicted score. In the quark sector, the $M_{11}$ matrix is approximately symmetric and can be block-diagonalized into $2\times 2$ blocks, each corresponding to a specific quark species. This confusion matrix provides a comprehensive overview of the model’s classification performance, highlighting both correct and incorrect predictions across various jet categories. These two state-of-the-art deep learning models show approximately similar performance as that of the BBT-Neutron. To further assess JoI performance, two key metrics are analyzed: jet flavor tagging efficiencies and charge flip rates. Jet flavor tagging efficiency is defined as half the sum of the values within each block, which does not distinguish the jet originating from quark and anti-quark. The charge flip rate is the ratio of the off-diagonal elements to the total sum of the block, which represents the probability of misidentification of the jet originating from quark and anti-quark. Figure 3 presents the flavor tagging efficiency and charge flip rates for each quark species, demonstrating that ParticleNet, Particle Transformer, and BBT-Neutron exhibit comparable performance at 10 million statistic.
@@
All three models show significant performance increase to dataset size scaling, forming an S-curve pattern in both flavor tagging efficiency and jet charge flip rates, as shown in Figures 4 and 5, respectively. Notably, distinct scaling behaviors were observed between BBT-Neutron and the specialized models, ParticleNet and Particle Transformer. The critical data thresholds on the S-curves for flavor tagging and charge flips suggest the emergence behaviors in BBT-Neutron, which were not observed in ParticleNet or Particle Transformer.
@@
In flavor tagging, ParticleNet and Particle Transformer, both specialized models, exhibit similar performance and scaling trends relative to the training data. This similarity is due to key structural and operational principles shared by both models, as described in Methods section 4.2. At smaller dataset sizes, these specialized models produce meaningful results, whereas BBT-Neutron initially performs at levels akin to random guessing. As the dataset size increases, all models exhibit performance gains, with the specialized models showing smooth improvement before reaching their respective performance plateaus. In contrast, BBT-Neutron requires larger datasets to achieve similar levels of performance, leading to a sharper jump on the curve. Particularly for b, c, and s quarks, BBT-Neutron requires approximately an order of magnitude more data than specialized models to achieve comparable performance. Once BBT-Neutron surpasses a critical dataset size threshold, its performance converges with that of the specialized models. For b,s and u quarks, BBT-Neutron exhibits a trend of surpassing the performance of the specialized models with further training on larger datasets.
@@
In the task of jet charge measurement a binary classification problem aimed at distinguishing jets originating from quarks of the same flavor but opposite charge—BBT-Neutron initially demonstrates near-zero performance, followed by a sharp improvement to above-random levels. This behavior is reminiscent of the emergent capabilities observed in large language models[40]. Specifically, BBT-Neutron produces random predictions when trained on datasets containing fewer than 10,000 events. However, its performance improves markedly beyond this threshold, suggesting that the model begins to capture complex, high-dimensional features after surpassing a critical data size. For datasets exceeding 30,000 events, BBT-Neutron’s scaling behavior aligns more closely with that of specialized models, exhibiting smoother performance gains with further increases in data size. By the time the dataset reaches approximately 3,000,000 events, BBT-Neutron achieves jet charge flip rates comparable to those of domain-specific models such as ParticleNet and Particle Transformer, which display steady performance improvements across all dataset sizes.
@@
The observed sharp jump in performance, from near-random to above-random levels, with increasing training dataset size in both flavor tagging and charge flip tasks for BBT-Neutron, strongly indicates the emergence of a generalist architecture. A plausible explanation for the absence of such emergent behaviors in specialist architectures, such as ParticleNet and Particle Transformer, is the incorporation of domain-specific structural biases. Both ParticleNet and Particle Transformer employ specialized architectural designs tailored to represent particle interactions and classification, which may lead to faster saturation of performance as data scales. In contrast, the generalist BBT-Neutron architecture processes all physical structures uniformly using a common data representation. Unlike ParticleNet and Particle Transformer, which leverage permutative invariance—a key feature aligned with fundamental symmetry laws—by eliminating positional encodings or related operations, BBT-Neutron employs a left-to-right sequence input, consistent with the sequence-to-sequence paradigm of language models. While this approach requires larger datasets for the generalist architecture to infer equivalent structures, it enables a pronounced performance leap once a critical dataset threshold is surpassed. The results further suggest that BBT-Neutron is capable of learning spatial symmetries embedded in the data, even without explicitly incorporating permutative invariance as in specialized models. Moreover, akin to the continuous phase transitions observed in the capabilities of large language models with increasing model size and computational scale, it is expected that generalist architectures like BBT-Neutron will exhibit further performance enhancements with continued training and scaling. In contrast to specialized architectures, LLMs like BBT-Neutron represent a paradigm shift in addressing scientific problems. BBT-Neutron exemplifies the potential of task-agnostic architectures capable of handling diverse scientific tasks such as classification, clustering, and regression through natural language inputs with minimal modifications to its output layer. This architectural strategy offers several key advantages:
@@
  • Reduced Development Overhead: Bin-BBT eliminates the need for task-specific architectural design and domain-specific feature engineering.
@@
3 Discussion and Future Work
In this study, we apply BBT-Neutron–with a task-agnostic architecture designed to address challenges in scientific problems enriched with large-scale numerical datasets–to the complex task of Jet Origin Identification (JoI) in high-energy physics, demonstrating that it achieves performance comparable to specialized AI tools like ParticleNet and Particle Transformer. Although BBT-Neutron requires larger datasets to match the accuracy of these specialized models, it scales effectively and delivers competitive results. This work underscores the potential of large language models for scientific domains that demand precise numerical representation and processing. BBT-Neutron lays the groundwork for further exploration of LLM applications in high-energy physics and other fields requiring sophisticated numerical data handling. Its scalability supports task-agnostic meta-learning and positions it as a versatile tool for data-intensive scientific tasks across disciplines.
@@ 4.1.1 Model architecture
Overview BBT-Neutron architecture involves converting input sequences into high-dimensional vectors through byte tokenization and patch embedding. The resulting embeddings pass through multiple transformer decoder layers. The model incorporates Rotary Position Embeddings (RoPE)[45] to infuse positional information into the self-attention mechanism, which allows our method to capture relative positional information between numbers, while Particle Transformer[46] lacks position encoding. Each layer of the transformer decoder is endowed with causal self-attention capabilities, ensuring a sequential, left-to-right processing flow and allowing for the intricate interplay among tokens within patches. This mechanism allows the model to discern the subtleties of numerical relationships, thereby revealing the underlying physical laws. A feed-forward neural network enhances the model’s ability to understand context and apply nonlinear transformations across layers. By layering these components, the model adeptly uncovers intricate dependencies within extended sequences. Finally, the output is transformed via a linear layer and a softmax function into a probability distribution over a vocabulary of size 257$\times$×Patch Size, predicting the next byte patch.
Furthermore, the BBT-Neutron model is versatile and can perform classification and regression tasks, which are common in many scientific applications where the goal is not necessarily to generate new sequences, but to classify inputs or predict continuous values [47, 48]. For classification tasks, the output dimension of the model corresponds to the number of distinct classes, allowing it to predict the most likely class label for a given input. In the case of regression tasks, the model outputs a single value representing the predicted outcome, such as a measurement or a score. This flexibility makes BBT-Neutron a powerful tool for a wide range of scientific analyses, from identifying patterns in experimental data to forecasting outcomes in complex systems.For classification, the output dimension is the number of class labels, and for regression, the output dimension is 1.
BBT-Neutron includes two versions, patch and nonpatch, which can be switched based on specific requirements.
  • Patch Version is a Patch-Based Model. It incorporates patches to enhance the overall computational speed of the model. As shown in Figure 6, patches allow the model to process input sequences more efficiently by breaking them into smaller, manageable segments. The v1 model is designed to be computationally efficient, making it suitable for tasks where speed is a critical factor.
@@
Patch Version models necessitate a careful approach to managing attention (padding) masks within each decoder layer. Each token only attends to its preceding tokens in the sequence, necessitating the masking of subsequent tokens. The byte-level attention mask must be reshaped to correspond with the structure of the patches. Following this, the boundary patches are marked as true for masking. This meticulous process ensures that the attention mechanism respects the patch boundaries accurately, thereby effectively identifying and integrating the inter dependencies both within individual patches and across different patches.
LM Head The output dimension of the language model head layer in the patch-based model is defined as Patch Size $\times$ 257. Here, 257 represents the total number of byte IDs, which include the byte values from 0 to 255, plus the padding ID represented by 256. And Patch Size is the number of patches into which the text sequence is divided. This design allows the model to generate predictions for each patch independently, maintaining the efficiency and effectiveness of the patch-based approach. In addition, for autoregressive generation in language models, generating byte-level tokens from patch-level representations presents certain challenges. Since patch-level embeddings summarize information from all bytes within a patch, they may lack some of the fine-grained details necessary for accurate byte-level token generation. This can make it more difficult to ensure that the generated tokens precisely reflect the intended content, especially in contexts where exact byte-level accuracy is crucial. To address this, the model may require additional mechanisms or fine-tuning to enhance the fidelity of byte-level token generation from patch-level embeddings.

### v2

@@ Abstract
For decades, researchers have developed task-specific models to address scientific challenges across diverse disciplines. Recently, large language models (LLMs) have shown enormous capabilities in handling general tasks; however, these models encounter difficulties in addressing real-world scientific problems, particularly in domains involving large-scale numerical data analysis, such as experimental high energy physics. This limitation is primarily due to BPE tokenization’s inefficacy with numerical data. In this paper, we propose a task-agnostic architecture, BBT-Neutron, which employs a binary tokenization method to facilitate pretraining on a mixture of textual and large-scale numerical experimental data. We demonstrate the application of BBT-Neutron to Jet Origin Identification (JoI), a critical categorization challenge in high-energy physics that distinguishes jets originating from various quarks or gluons. Our results indicate that BBT-Neutron achieves comparable performance to state-of-the-art task-specific JoI models. Furthermore, we examine the scaling behavior of BBT-Neutron’s performance with increasing data volume, suggesting the potential for BBT-Neutron to serve as a foundational model for particle physics data analysis, with possible extensions to a broad spectrum of scientific computing applications for Big Science experiments, industrial manufacturing and spacial computing. The project code is available at https://github.com/supersymmetry-technologies/bbt-neutron.
@@ 1 Introduction
Scaling law plays a critical role in developing emergent abilities and in-context learning of large language models [1, 2, 3, 4]. Pretraining LLM on a sufficiently broad and heterogeneous corpus gives rise to in-context learning building upon the processes of absorbing skills within the repetitive sequences of subsets of data which usually come as natural language [5, 6]. The efficacy of in-context learning is directly proportional to the volume and variety of the data, as well as the computational power expended during the pretraining phase, adhering to a power-law distribution [7]. The “emergence” of LLM, described by the ability to solve specific tasks jumping up abruptly when data and model size pass a certain threshold, is attributed to the scaling of pretraining.
@@
First, the mainstream number representation method, Byte Pair Encoding (BPE) tokenization, as adopted by GPT-3 [5, 6], limits large language models’ (LLMs) performance in numerical processing and reasoning tasks. BPE tokenization, a simple data compression technique that iteratively replaces the most frequent byte pairs with a single, unused byte can introduce ambiguity when tokenizing numbers, as it often fragments numerical values into arbitrary segments, thereby obscuring their original meaning. Additionally, BPE’s statistical approach may result in inconsistent tokenization of identical numbers across different contexts, further complicating downstream numerical tasks. Llama3 [17] attempts to address this by splitting numbers into indi-vidual tokens to improve consistency in number representation. However, even the most advanced number tokenization techniques cannot eliminate the fundamental issue: the quantitative relationships between individual numbers are distorted post-tokenization, impairing foundational models’ ability to perform basic comparisons and arithmetic operations [9, 18, 19]. This limitation has driven consensus on the necessity of advancing number tokenization methods to enable significant improvements in LLM applications for scientific tasks.
@@
Third, foundational models have been applied to scientific tasks in areas such as question answering [23, 24], quantum physics calculations [25], prime number distribution [26], partial differential equation (PDE) computation [27], and time series regression [28]. However, to date, LLMs developed for scientific applications are primarily trained on symbolic data, including text and scientific symbols such as DNA sequences and mathematical formulas. Numerical experimental data particularly from large-scale scientific projects in fields like particle physics and astronomy are typically excluded from these training corpora due to the absence of a general architecture capable of multimodal synthesis of text and experimental data. This limitation significantly hinders the capacity of general LLMs to address complex, real-world scientific problems.
@@
JoI is a newly developed idea for experimental particle physics that aims to distinguish jets originating from different quarks and gluons, which are the building blocks of our material world and are fundamental particles in the Standard Model of Particle Physics. Once quark or gluon is produced in high-energy collisions, such as the Large Hadron Collider (LHC) and future electron-positron Higgs factories, it would immediately give rise to a spray of particles, mostly hadrons, that travel in the same direction. This spray of particles is conventionally known as a jet, which is a critical object for physics measurements in collider experiments. Significant efforts have been devoted to identifying the origin of jets, specifically, determining which type of colored particle initiated the jet. Identification of jet origin is central to many physics analyses, particularly in studies of the Higgs, W, and Z bosons, as nearly 70% of these bosons decay into two jets directly. In addition, jets are fundamental to advancing our understanding of QCD. However, jets originating from different types of colored particles exhibit only marginal differences in their observables, making accurately identifying the origin of the jet extremely challenging.
@@
JoI was incarnated in the specific algorithms of jet flavor tagging, jet charge measurement, and strange, light, and gluon jet tagging. At the Large Electron-Positron Collider (LEP), the predecessor to the LHC, jet origin identification initially relied on cuts applied to various discriminating variables to distinguish between b-, c-, and light-jets [30, 31]. Over time, likelihood-based methods [32, 33] emerged, combining multiple observables into a single discriminant. With advancements in detector technology, the LHC provided improved reconstruction performance, coupled with advanced algorithms such as Boosted Decision Trees (BDTs) [34, 35], followed by recurrent neural networks (RNNs) [36, 37], convolutional neural networks (CNNs) [38, 39], Graph Neural Network (GNN), and transformer, the origin of jet has been identified with better performance. JoI was then developed using GNN-based ParticleNet and Transfomer-based Particle Transformer; it could simultaneously identify 11 jet species: five quarks, five anti-quarks, and gluons at the proposed electron-positron Higgs factory. JoI offers an unprecedented distinguishing power of different jet species and is critical for almost all the physics measurements with the jet’s final state. Quantitative analysis shows that the anticipated accuracies of relevant Higgs measurements could be improved from two folds up to two orders of magnitude using JoI, and thus significantly boost the discovery power of collider experiments.
@@
We observe that BBT-Neutron achieves a performance comparable to that of specialized models like ParticleNet and Particle Transformer. As the first task-agnostic general LLM architecture capable of pretraining on a multimodal mixture of textual and large-scale numerical datasets, BBT-Neutron demonstrates state of the art performance in JoI, compared to the results of leading domain-specific models. This achievement marks a significant step toward developing foundational models for scientific research applications that can generalize across diverse data types and tasks.
@@ 2 Results
We compare the performance of BBT-Neutron and two state-of-the-art models, ParticleNet(PN) [50] and Particle Transformer(ParT) [46]. The results for these models are shown in Fig 2, using an 11-dimensional confusion matrix $M_{11}$, which classifies each jet according to the category with the highest predicted score. In the quark sector, the $M_{11}$ matrix is approximately symmetric and can be block-diagonalized into $2\times 2$ blocks, each corresponding to a specific quark species. This confusion matrix provides a comprehensive overview of the model’s classification performance, highlighting both correct and incorrect predictions across various jet categories. The top panel of Fig 2 presents the overall JoI performance for BBT-Neutron. These two state-of-the-art deep learning models show approximately similar performance as that of the BBT-Neutron. To further assess JoI performance, two key metrics are analyzed: jet flavor tagging efficiencies and charge flip rates. Jet flavor tagging efficiency is defined as half the sum of the values within each block, which does not distinguish the jet originating from quark and anti-quark. The charge flip rate is the ratio of the off-diagonal elements to the total sum of the block, which represents the probability of misidentification of the jet originating from quark and anti-quark. Figure 3 presents the flavor tagging efficiency and charge flip rates for each quark species, demonstrating that PN, ParT, and BBT-Neutron exhibit comparable performance at 10 million statistic.
@@
We conduct a comparative analysis on the scaling behaviors of BBT-Neutron and two state-of-the-art models, PN and ParT, applied to JOI across training datasets ranging from 100 to 10 million events. All three models show significant performance increase to dataset size scaling, forming an S-curve pattern in both flavor tagging efficiency and jet charge flip rates, as shown in Figures 4 and 5, respectively. Notably, distinct scaling behaviors were observed between BBT-Neutron and the specialized models. The critical data thresholds on the S-curves for flavor tagging and charge flips indicate distinct emergence behaviors in BBT-Neutron. These behaviors are not as pronounced in the specialized models. In flavor tagging, PN and ParT, both specialized models, exhibit similar performance and scaling trends relative to the training data. This similarity is due to key structural and operational principles shared by both models, as described in Methods section 4.2. At smaller dataset sizes, these specialized models produce meaningful results, whereas BBT-Neutron initially performs at levels akin to random guessing. As the dataset size increases, all models exhibit performance gains, with the specialized models showing smooth improvement before reaching their respective performance plateaus. In contrast, BBT-Neutron requires larger datasets to achieve similar levels of performance, leading to a sharper jump on the curve. Particularly for b, c, and s quarks, BBT-Neutron requires approximately an order of magnitude more data than specialized models to achieve comparable performance. Once BBT-Neutron surpasses a critical dataset size threshold, its performance converges with that of the specialized models. For b,s and u quarks, BBT-Neutron exhibits a trend of surpassing the performance of the specialized models with further training on larger datasets.
@@
In the task of jet charge measurement a binary classification problem aimed at distinguishing jets originating from quarks of the same flavor but opposite charge. BBT-Neutron initially demonstrates near-random performance, followed by a sharp improvement to above-random levels. This behavior is reminiscent of the emergent capabilities observed in large language models [40]. Specifically, BBT-Neutron produces random predictions when trained on datasets containing fewer than 10,000 events. However, its performance improves markedly beyond this threshold, suggesting that the model begins to capture complex, high-dimensional features after surpassing a critical data size. For datasets exceeding 30,000 events, BBT-Neutron’s scaling behavior aligns more closely with that of specialized models, exhibiting smoother performance gains with further increases in data size. By the time the dataset reaches approximately 3,000,000 events, BBT-Neutron achieves jet charge flip rates comparable to those of domain-specific models, which display steady performance improvements across all dataset sizes.
@@
The observed sharp jump in performance, from near-random to above-random levels, with increasing training dataset size in both flavor tagging and charge flip tasks for BBT-Neutron, strongly indicates the emergence of a generalist architecture. A plausible explanation for the absence of such emergent behaviors in specialist architectures, such as PN and ParT, is the incorporation of domain-specific structural biases. These models employ specialized architectural designs tailored to represent particle interactions and classification, which may lead to faster saturation of performance as data scales. In contrast, the generalist BBT-Neutron architecture processes all physical structures uniformly using a common data representation. Unlike specialist architectures, which leverage permutative invariance-a key feature aligned with fundamental symmetry laws-by eliminating positional encodings or related operations, BBT-Neutron employs a left-to-right sequence input, consistent with the sequence-to-sequence paradigm of language models. While this approach requires larger datasets for the generalist architecture to infer equivalent structures, it enables a pronounced performance leap once a critical dataset threshold is surpassed. The results further suggest that BBT-Neutron is capable of learning spatial symmetries embedded in the data, even without explicitly incorporating permutative invariance as in specialized models.
3 Discussion and Future Work
In this study, we apply BBT-Neutron with a task-agnostic architecture designed to address challenges in scientific problems enriched with large-scale numerical datasets to the complex task of Jet Origin Identification (JoI) in high-energy physics, demonstrating that it achieves performance comparable to specialized AI tools like PN and ParT. Although BBT-Neutron requires larger datasets to match the accuracy of these specialized models, it scales effectively and delivers competitive results. This work underscores the potential of large language models for scientific domains that demand precise numerical representation and processing. BBT-Neutron lays the groundwork for further exploration of LLM applications in high-energy physics and other fields requiring sophisticated numerical data handling. Its scalability supports task-agnostic meta-learning and positions it as a versatile tool for data-intensive scientific tasks across disciplines.
Moreover, akin to the continuous phase transitions observed in the capabilities of large language models with increasing model size and computational scale, it is expected that generalist architectures like BBT-Neutron will exhibit further performance enhancements with continued training and scaling. In contrast to specialized architectures, LLMs like BBT-Neutron represent a paradigm shift in addressing scientific problems. BBT-Neutron exemplifies the potential of task-agnostic architectures capable of handling diverse scientific tasks such as classification, clustering, and regression through natural language inputs with minimal modifications to its output layer. This architectural strategy offers several key advantages:
@@ 3 Discussion and Future Work
  • Reduced Development Overhead: BBT-Neutron eliminates the need for task-specific architectural design and domain-specific feature engineering.
@@ 4.1.1 Model architecture
Overview BBT-Neutron architecture involves converting input sequences into high-dimensional vectors through byte tokenization and patch embedding. The resulting embeddings pass through multiple transformer decoder layers. The model incorporates Rotary Position Embeddings (RoPE) [45] to infuse positional information into the self-attention mechanism, which allows our method to capture relative positional information between numbers, while Particle Transformer [46] lacks position encoding. Each layer of the transformer decoder is endowed with causal self-attention capabilities, ensuring a sequential, left-to-right processing flow and allowing for the intricate interplay among tokens within patches. This mechanism allows the model to discern the subtleties of numerical relationships, thereby revealing the underlying physical laws. A feed-forward neural network enhances the model’s ability to understand context and apply nonlinear transformations across layers. By layering these components, the model adeptly uncovers intricate dependencies within extended sequences. Finally, the output is transformed via a linear layer and a softmax function into a probability distribution over a vocabulary of size Patch Size×$\times$257, predicting the next byte patch.
Furthermore, the BBT-Neutron model is versatile and can perform classification and regression tasks, which are common in many scientific applications where the goal is not necessarily to generate new sequences, but to classify inputs or predict continuous values [47, 48]. For classification tasks, the output dimension of the model corresponds to the number of distinct classes, allowing it to predict the most likely class label for a given input. In the case of regression tasks, the model outputs a single value representing the predicted outcome, such as a measurement or a score. This flexibility makes BBT-Neutron a powerful tool for a wide range of scientific analyses, from identifying patterns in experimental data to forecasting outcomes in complex systems.
BBT-Neutron includes two versions, patch and non-patch, which can be switched based on specific requirements.
  • Patch Version is a Patch-Based Model. It incorporates patches to enhance the overall computational speed of the model. As shown in Figure 6, patches allow the model to process input sequences more efficiently by breaking them into smaller, manageable segments. The Patch Version model is designed to be computationally efficient, making it suitable for tasks where speed is a critical factor.
@@
Patch Version models necessitate a careful approach to managing attention masks within each decoder layer. Attention Mask is used to inform the model which parts of the input are actual data and which parts are padding, the model needs to disregard the padded parts because the padding tokens do not contain semantic information. The attention mask must be reshaped to correspond with shape of the patches. Following this, the boundary patches are marked as true for masking. This meticulous process ensures that the attention mechanism respects the patch boundaries accurately, thereby effectively identifying and integrating the inter dependencies both within individual patches and across different patches.
LM Head The output dimension of the language model head layer in the patch-based model is defined as Patch Size $\times$ 257. Here, 257 represents the total number of byte IDs, which include the byte values from 0 to 255, plus the padding ID represented by 256, and Patch Size is the number of patches into which the text sequence is divided. This design allows the model to generate predictions for each patch independently, maintaining the efficiency and effectiveness of the patch-based approach. In addition, for autoregressive generation in language models, generating byte-level tokens from patch-level representations presents certain challenges. Since patch-level embeddings summarize information from all bytes within a patch, they may lack some of the fine-grained details necessary for accurate byte-level token generation. This can make it more difficult to ensure that the generated tokens precisely reflect the intended content, especially in contexts where exact byte-level accuracy is crucial. To address this, the model may require additional mechanisms or fine-tuning to enhance the fidelity of byte-level token generation from patch-level embeddings.